# COSMOS: Cross-Modality Self-Distillation for Vision Language Pre-training

### v1

@@ Abstract
Vision-Language Models (VLMs) trained with contrastive loss have achieved significant advancements in various vision and language tasks. However, the global nature of contrastive loss makes VLMs focus predominantly on foreground objects, neglecting other crucial information in the image, which limits their effectiveness in downstream tasks. To address these challenges, we propose COSMOS: CrOSs-MOdality Self-distillation for vision-language pre-training that integrates a novel text-cropping strategy and cross-attention module into a self-supervised learning framework. We create global and local views of images and texts (i.e., multi-modal augmentations), which are essential for self-distillation in VLMs. We further introduce a cross-attention module, enabling COSMOS to learn comprehensive cross-modal representations optimized via a cross-modality self-distillation loss. COSMOS consistently outperforms previous strong baselines on various zero-shot downstream tasks, including retrieval, classification, and semantic segmentation. Additionally, it surpasses CLIP-based models trained on larger datasets in visual perception and contextual understanding tasks. Code and models will be open-sourced at ExplainableML/cosmos.
@@ 1 Introduction
Previous works [43, 52, 53, 21, 61] have tackled these limitations by masking the text or the image [21, 43, 61], or explicitly optimizing local-to-global image representation. For instance, SLIP [52] and SILC [53] improved the visual representations by integrating self-supervision into the CLIP image encoder. However, we argue that enhancing only the image encoder leads to sub-optimal results, since this does not fully exploit the multi-modality of the data. Therefore, we propose COSMOS, to learn CrOSs-MOdality embeddings through Self-distillation. Unlike previous self-supervised methods [52, 53, 21], COSMOS incorporates a unique text-cropping strategy and a cross-attention module to learn multi-modal representations for various downstream tasks.
@@
Our text-cropping strategy (Fig. 1, top) is inspired by the multi-crop augmentation for images [7] which optimizes local-to-global image correspondence during training. We randomly sample from the long synthetic caption datasets [83, 65] to construct sentences of varying length. By applying multi-crop augmentation to both images and text, we enhance the visual and text encoders simultaneously within a self-distillation framework [7], aligning the representation of the student VLM with that of the teacher.
@@
Our cross-attention module integrated in the student allows the model to attend to the other modality’s representation, ensuring cross-modality information is captured during training. Inside the module, image patch tokens are conditioned on texts or word tokens are conditioned on images to learn visual and textual grounding. As illustrated in Fig. 1 (bottom), our cross-attention module focuses on different objects from the entire image, which are highlighted in both images and captions based on the attention weights. This demonstrates that COSMOS is able to effectively localize relevant information in images and captions.
@@
We demonstrate the adaptability of COSMOS by pre-training it on four datasets varying from 3M to 30M pairs. For evaluating our method, we perform zero-shot experiments on 2 retrieval, 11 classification, and 6 segmentation benchmarks, as well as, the visual perception [69] and contextual understanding tasks [33, 32]. The extensive experiments on these downstream tasks reveal that our method consistently outperforms the previous CLIP-based methods, under the same pre-training data budget. COSMOS pre-trained on only 12 million examples outperforms the OpenCLIP [13] models trained on billions of image-text pairs, in both image-to-text and text-to-image retrieval tasks.
@@
Our contributions can be summarized as follows: (1) We design a text-cropping strategy which promotes local-to-global text representation learning, fully leveraging multi-modality in the self-distillation of VLMs. (2) We demonstrate that cross-modality embeddings are effective to simultaneously self-distill both image and text encoders. (3) COSMOS reveals fine-grained perception of images and compositional understanding of language, outperforming baselines trained on larger datasets.
@@ 2 Related Work
Figure 2: An overview of COSMOS. Left: Our VLM pre-training mechanism is based on student-teacher framework with contrastive loss ($\mathcal{L}_{\text{CLIP}}$) for multi-modal alignment and cross-modality self-distillation loss ($\mathcal{L}_{\text{COSMOS}}$) for fine-grained representation learning. Right: The architecture of the student and teacher model with cross-attention modules that extracts cross-modality information from the student.
@@
Self-Supervised Learning (SSL) learns discriminative features from unlabeled data [30]. Initially, models like BERT [20] and GPT [58] were designed to predict masked words or the next word in a sequence, eliminating the need for additional labels. Similarly, vision models benefit from self-supervised techniques, either with contrastive learning or masked image modeling. Based on contrastive learning, BYOL [29] and DINO [7] trained a student network to produce image embeddings that closely resemble those of a teacher network, with the teacher being an exponential moving average of the student. On the other hand, masked image modeling divided the input image into visual tokens and predicted a randomly masked subset of them. For example, MAE [31] directly predicted the original pixel from the masked patches, and BEiT [3] employed a pre-trained tokenizer as its target.
@@
Contrastive learning in vision-language pre-training has achieved great success in various downstream tasks, demonstrating strong zero-shot transfer ability [59, 34, 77, 26, 28, 40, 27]. As these models are trained to learn visual representations aligned with the language representation, most MLLMs [1, 41, 47, 46, 86] adopt these vision encoders to process the image input. However, instance-level contrastive learning suffers from feature suppression, where the model learns only the dominant features in the data while neglecting other valuable features [12, 60, 82, 42, 66, 75, 44, 69]. In other words, the model creates so-called simple shortcut features and decision rules that do not consider all available distinguishing features. Some works [10, 65, 70, 81, 55, 83, 72] have employed datasets equipped with long synthetic captions generated by MLLMs. These datasets help mitigate feature suppression by learning different features described in the captions and avoiding overfitting on a single caption. Referring to Zheng et al. [83], we employ image datasets with long synthetic captions. Instead of simply training our model with these additional captions, we create text augmentation with the synthetic captions and introduce the cross-attention module for self-distillation.
@@
Following the success of SSL, recent works [43, 52, 53, 21, 61, 36] have applied SSL objectives to contrastive vision-language pre-training. These approaches also tackled the feature suppression limitation, where the new objective reduced the inductive bias to prevent the creation of shortcut features. For example, SLIP [52] combined the SimCLR [11] loss with the contrastive loss, resulting in better transferability compared to the original CLIP. SILC [53] explicitly matched local crops to global crops during training inspired by DINO [7], while MaskCLIP [21] learned more fine-grained features by applying masked modeling to both images and text. Similar to SLIP [52] and SILC [53], we also employ self-supervised learning for vision-language pre-training. While SLIP [52] and SILC [53] only improved the image representation without improving the text representation, we propose a text-cropping technique that enables us to improve the text and image representations simultaneously with our novel cross-modality self-distillation loss.
@@ 3 COSMOS Model
We describe our augmentation technique, which is the key component of self-distillation in Sec. 3.1, COSMOS’s architecture in Sec. 3.2, and the training objective in Sec. 3.3.
@@ 3.1 Text and Image Augmentations
Text Cropping Strategy. For performing self-supervised learning based on text, we propose a text-cropping strategy utilizing long caption datasets [83, 65]. These datasets consist of images paired with detailed synthetic captions generated by MLLMs. Inspired by the multi-crop strategy for images [7], we define global and local concepts for text captions. The global crop of a caption comprises one to five sentences randomly sampled from the long synthetic caption, while the local crop of a caption consists of a single sentence, also randomly sampled from the long caption. The number of sentences in the global captions is randomly determined at each training iteration. This adaptation ensures that the global caption generally describes larger areas of the original image, e.g. the fox with vibrant red fur is the main subject of the image, in Fig. 1 top, while the local caption focuses on relatively smaller regions, e.g. the snow covers the ground and the trees.
@@ 3.3 Training Objective
The total contrastive loss $\mathcal{L}_{\text{CLIP}}$ is calculated between the image and text embeddings of the student as follows:
@@
As mentioned in Sec. 3.1, all crops are passed through the student while only global crops are given to the teacher. Thus, $\mathcal{L}_{\text{CLIP}}$ is calculated between all crops within the student, optimizing the model to construct the overall structure of image and text representations by aligning the embeddings through contrastive learning. On the other hand, $\mathcal{L}_{\text{COSMOS}}$ is computed between global crops given to the teacher and all crops given to the student, encouraging the model to learn fine-grained representations by predicting the rich global context of the teacher from the local feature of the student (the detailed pseudocode of the training procedure is in supplementary Sec. C.3).
@@ 4 Experiments
Table 2: Zero-shot classification results in terms of top-1 accuracy on ImageNet [18] and Food101 [4], CIFAR-10 [38], CIFAR-100 [38], SUN397 [73], Stanford Cars [37], FGVCAircraft [50], DTD [14], Oxford Pets [57], Caltech101 [24], Flowers102 [54], and ImageNet [18] datasets. The vision encoder architecture is ViT-B/16. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
Pre-training Datasets. We pre-train COSMOS on Conceptual Caption 3M (CC3M) and 12M (CC12M) [64], as well as, YFCC15M [16] a subset of YFCC100M [67]. The combination of these three datasets (CC3M, CC12M, and YFCC15M) results in a merged dataset of 30M pairs, referred to as Merged-30M, following DreamLIP [83]. As mentioned in Sec. 3.1, we utilize synthetic long captions to obtain multi-modal augmentation. For CC3M, CC12M and YFCC15M, we use the synthetic captions generated by Zheng et al. [83]. The new captions were obtained using InstructBLIP [17], LLaVA-1.5 [46], and ShareGPT4V [10].
Implementation Details. COSMOS is built on the OpenCLIP [13] code repository, where we primarily adopt the default settings of their implementation. We use ViT-B/16 as the vision encoder designed for $224\times 224$ image sizes, while our CLIP text encoder has the sequence length of 77. Each encoder consists of 12 transformer layers, and the embedding size is fixed at 512. We train COSMOS on A100 GPUs for 32 epochs, using a learning rate of $5\times 10^{-4}$ and a batch size of 1,024 for CC3M, or 4,096 for CC12M, YFCC15M and Merged-30M. Following DINO [7], we crop two global image views at scales between 0.4 and 1.0, resizing them to $224\times 224$ pixels, and six local image views at scales between 0.05 and 0.4, resized to $96\times 96$ pixels. Similarly, the two global text views consist of one to five sentences randomly sampled from the synthetic captions, while the six local text views consist of a single sentence sampled from the same caption set. The teacher momentum $\lambda$ is empirically set to 0.999 for CC3M and 0.99 for other datasets (more details in supplementary Sec. C.1).
@@ 4.1 Zero-Shot Image-Text Retrieval
As shown in Tab. 1, COSMOS achieves 53.1% and 40.1% R@1 score for image-to-text (I2T) and text-to-image (T2I) retrieval on MSCOCO [45]. These results are 12.3% and 10.3% higher than DreamLIP [83] (40.8% and 29.8%, respectively). When we switch to a larger pre-training set (CC12M), COSMOS further improves its performance, reaching 64.2% and 48.9% R@1 scores, outperforming CLIP [59] by 8.0% and 9.1%, respectively. When we employ YFCC15M [16] as the pre-training dataset, COSMOS attains a performance of 67.5% and 50.9% in terms of I2T and T2I R@1 scores. The highest performance of 68.0% and 52.5% on the MSCOCO [45] dataset is obtained by COSMOS, when Merged-30M is used for pre-training. COSMOS also achieves the best performance for image-text retrieval on the Flickr30k [78] dataset, achieving an R@1 score of 92.9% and 80.3% in terms of I2T and T2I retrieval when Merged-30M is used for pre-training, while SigLIP [80] attains only 89.9% and 73.6%, respectively. Overall, COSMOS outperforms all baselines on both datasets (Flickr30k [78] and MSCOCO [45]) in the image-text retrieval task.
@@
Table 3: Zero-shot semantic segmentation results in terms of mean Intersection over Union (mIoU) on PASCAL VOC 2012 [22], PASCAL Context [51], Cityscapes [15], ADE20k [84], COCO-Stuff, and COCO-Object [5] with and without the background category. The vision encoder architecture is ViT-B/16. We also report the performance of our method enhanced with the SCLIP mechanism [71]. The best results are highlighted in bold. : Results copied from their work. : Results obtained using their official pre-trained weights.
@@ 4.2 Zero-Shot Image Classification
We observe in Tab. 2 that COSMOS outperforms the baselines CLIP [59], SigLIP [80], DreamLIP [83], MaskCLIP [21], and SF-CLIP [61] on most of the evaluation datasets, when the pre-training dataset contains the same number of samples. For instance, COSMOS reaches an accuracy of 37.1% on ImageNet when the pre-training dataset is CC3M. As expected, the classification task greatly benefits from the pre-training dataset size, i.e. COSMOS pre-trained on CC12M achieves a score of 51.4% on ImageNet, reaching 57.6% when pre-trained on Merged-30M.
Additionally, COSMOS surpasses all baselines regardless of the pre-training set, in terms of average accuracy computed over the 11 datasets. For instance, when pre-trained on YFCC15M, COSMOS achieves an average accuracy of 50.0%, while SigLIP [80] and DreamLIP [83] reach 44.2% and 46.3%, respectively. It is interesting to note that on two fine-grained datasets, i.e. Cars and SUN397, the accuracy improvements especially for Merged-30M trained COSMOS is significant (67.8% vs 64.7% on SUN397, 31.2% vs 23.4% on Cars, compared to DreamLIP [83]).
For reference, models trained on billions of samples, e.g. OpenCLIP [13], MetaCLIP [74], and Llip [39] achieve higher accuracy, implying that the larger the number of images the higher the accuracy for the classification task, since the model is required to learn the global information.

### v2

@@ Abstract
Vision-Language Models (VLMs) trained with contrastive loss have achieved significant advancements in various vision and language tasks. However, the global nature of the contrastive loss makes VLMs focus predominantly on foreground objects, neglecting other crucial information in the image, which limits their effectiveness in downstream tasks. To address these challenges, we propose COSMOS: CrOSs-MOdality Self-distillation for vision-language pre-training that integrates a novel text-cropping strategy and cross-attention module into a self-supervised learning framework. We create global and local views of images and texts (i.e., multi-modal augmentations), which are essential for self-distillation in VLMs. We further introduce a cross-attention module, enabling COSMOS to learn comprehensive cross-modal representations optimized via a cross-modality self-distillation loss. COSMOS consistently outperforms previous strong baselines on various zero-shot downstream tasks, including retrieval, classification, and semantic segmentation. Additionally, it surpasses CLIP-based models trained on larger datasets in visual perception and contextual understanding tasks. Code is available at https://github.com/ExplainableML/cosmos.
@@ 1 Introduction
Previous works [45, 56, 57, 21, 65] have tackled these limitations by masking the text or the image [21, 45, 65] and explicitly optimizing local-to-global image representation. For instance, SLIP [56] and SILC [57] improved the visual representations by integrating self-supervision into the CLIP image encoder. However, we argue that enhancing only the image encoder leads to sub-optimal results, since this does not fully exploit the multi-modality of the data. Therefore, we propose COSMOS, to learn CrOSs-MOdality embeddings through Self-distillation. Unlike previous self-supervised methods [56, 57, 21], COSMOS incorporates a unique text-cropping strategy and a cross-attention module to learn multi-modal representations for various downstream tasks.
@@
Our text-cropping strategy (Fig. 1, top) is inspired by the multi-crop augmentation for images [7], which optimizes local-to-global image correspondence during training. We randomly sample from the long synthetic caption datasets [91, 70] to construct sentences of varying length. By applying multi-crop augmentation to both images and text, we enhance the visual and text encoders simultaneously within a self-distillation framework [7], aligning the representation of the student VLM with that of the teacher.
@@
Our cross-attention module, integrated into the student, allows the model to attend to the other modality’s representation, ensuring that cross-modality information is captured during training. Inside the module, image patch tokens are conditioned on texts, or word tokens are conditioned on images, to learn visual and textual grounding. As illustrated in Fig. 1 (bottom), our cross-attention module focuses on different objects from the entire image, which are highlighted in both images and captions based on the attention weights. This demonstrates that COSMOS is able to effectively localize relevant information in images and captions.
@@
We demonstrate the adaptability of COSMOS by pre-training it on 4 datasets varying from 3M to 30M pairs. To evaluate our method, we perform zero-shot experiments on 2 retrieval, 11 classification, and 6 segmentation benchmarks, as well as, the visual perception [74] and contextual understanding tasks [33, 32]. The extensive experiments on these downstream tasks reveal that our method consistently outperforms the previous CLIP-based methods, under the same pre-training data budget. Moreover, COSMOS pre-trained on only 12 million examples outperforms the OpenCLIP [13] models trained on billions of image-text pairs, in both image-to-text and text-to-image retrieval tasks.
@@
Our contributions are summarized as follows: (1) We design a text-cropping strategy which promotes local-to-global text representation learning, fully leveraging multi-modality in the self-distillation of VLMs. (2) We demonstrate that cross-modality embeddings are effective to simultaneously self-distill both image and text encoders. (3) COSMOS reveals fine-grained perception of images and compositional understanding of language, outperforming baselines trained on larger datasets.
@@ 2 Related Work
Figure 2: An overview of COSMOS. Left: Our VLM pre-training mechanism is based on the student-teacher framework with contrastive loss ($\mathcal{L}_{\text{CLIP}}$) for multi-modal alignment, and cross-modality self-distillation loss ($\mathcal{L}_{\text{COSMOS}}$) for fine-grained representation learning. Right: The architecture of the student and teacher model with cross-attention modules that extract cross-modality information from the student.
@@
Self-Supervised Learning (SSL) learns discriminative features from unlabeled data [30]. Initially, models like BERT [20] and GPT [62] were designed to predict masked words or the next word in a sequence, eliminating the need for additional labels. Similarly, vision models benefit from self-supervised techniques, either with contrastive learning or masked image modeling. Based on contrastive learning, BYOL [29] and DINO [7] trained a student network to produce image embeddings that closely resemble those of a teacher network, with the teacher being an exponential moving average of the student. On the other hand, masked image modeling divides the input image into visual tokens and predicts a randomly masked subset of them. For example, MAE [31] directly predicted the original pixel from the masked patches, and BEiT [3] employed a pre-trained tokenizer as its target.
@@
Contrastive learning in vision-language pre-training has achieved great success in various downstream tasks, demonstrating strong zero-shot transfer abilities [63, 35, 83, 26, 28, 42, 27]. As these models are trained to learn visual representations aligned with the language representation, most Multi-modal Large Language Models (MLLMs) [1, 43, 50, 49, 94] adopt these vision encoders to process the image input. However, instance-level contrastive learning suffers from feature suppression, where the model learns only the dominant features in the data while neglecting other valuable features [12, 64, 90, 44, 71, 81, 47, 74]. In other words, the model creates so-called simple shortcut features and decision rules that do not consider all available distinguishing features. Some works [10, 70, 75, 89, 59, 91, 78] have employed datasets equipped with long synthetic captions generated by MLLMs. These datasets help mitigate feature suppression by learning different features described in the captions and avoiding overfitting to a single caption. Referring to Zheng et al. [91], we employ image datasets with long synthetic captions. Instead of simply training our model with these additional captions, we create text augmentation using the synthetic captions, and introduce the cross-attention module for self-distillation.
@@
Following the success of SSL, recent works [45, 56, 57, 21, 65, 37] apply SSL objectives to contrastive vision-language pre-training. These approaches also tackle the feature suppression limitation, where the new objective reduces the inductive bias to prevent the creation of shortcut features. For example, SLIP [56] combined the SimCLR [11] loss with the contrastive loss, resulting in better transferability compared to the original CLIP. Similarly, SILC [57] explicitly matched local crops to global crops during training inspired by DINO [7], while MaskCLIP [21] learned more fine-grained features by applying masked modeling to both images and text. Similar to SLIP [56] and SILC [57], we also employ self-supervised learning for vision-language pre-training. While SLIP [56] and SILC [57] only improve the image representation without improving the text representation, we propose a text-cropping technique that enables us to improve the text and image representations simultaneously with our novel cross-modality self-distillation loss.
@@ 3 COSMOS Model
We describe our augmentation technique, the key component in self-distillation, in Sec. 3.1, COSMOS’s architecture in Sec. 3.2, and the training objective in Sec. 3.3.
@@ 3.1 Text and Image Augmentations
Text Cropping Strategy. To perform self-supervised learning based on text, we propose a text-cropping strategy utilizing long caption datasets [91, 70]. These datasets consist of images paired with detailed synthetic captions generated by MLLMs. Inspired by the multi-crop strategy for images [7], we define global and local concepts for text captions. The global crop of a caption comprises one to five sentences randomly sampled from the long synthetic caption, while the local crop of a caption consists of a single sentence, also randomly sampled from the long caption. The number of sentences in the global captions is randomly determined at each training iteration. This adaptation ensures that the global caption generally describes larger areas of the original image, e.g. the fox with vibrant red fur is the main subject of the image, in Fig. 1 top, while the local caption focuses on relatively smaller regions, e.g. the snow covers the ground and the trees.
@@ 3.3 Training Objective
The total contrastive loss $\mathcal{L}_{\text{CLIP}}$ is calculated between the image and text embeddings of the student as follows.
@@
As mentioned in Sec. 3.1, all crops are passed through the student while only global crops are given to the teacher. Thus, $\mathcal{L}_{\text{CLIP}}$ is calculated between all crops within the student, optimizing the model to construct the overall structure of image and text representations by aligning the embeddings through contrastive learning. On the other hand, $\mathcal{L}_{\text{COSMOS}}$ is obtained between global crops given to the teacher and all crops given to the student, encouraging the model to learn fine-grained representations by predicting the rich global context of the teacher from the local feature of the student (the detailed pseudocode of the training objective is in Supplementary Material Algorithm 1).
@@ 4 Experiments
Table 2: Zero-shot classification results in terms of top-1 accuracy on ImageNet [18] and Food101 [4], CIFAR-10 [39], CIFAR-100 [39], SUN397 [79], Stanford Cars [38], FGVCAircraft [54], DTD [14], Oxford Pets [61], Caltech101 [24], Flowers102 [58], and ImageNet [18]. The vision encoder architecture is ViT-B/16. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
Pre-training Datasets. We pre-train COSMOS on Conceptual Caption 3M (CC3M) and 12M (CC12M) [68] as well as YFCC15M [16], a subset of YFCC100M [72]. The combination of these three datasets (CC3M, CC12M, and YFCC15M) results in a merged dataset of 30M pairs, referred to as Merged-30M, following DreamLIP [91]. As mentioned in Sec. 3.1, we utilize synthetic long captions to obtain multi-modal augmentation. For CC3M, CC12M and YFCC15M, we use the synthetic captions generated by DreamLIP [91]. The new captions were obtained using InstructBLIP [17], LLaVA-1.5 [49], and ShareGPT4V [10].
Implementation Details. COSMOS is built on the OpenCLIP [13] code repository, where we primarily adopt the default settings of their implementation. We use ViT-B/16 as the vision encoder designed for $224\times 224$ image sizes, while our CLIP text encoder has the sequence length of 77. Each encoder consists of 12 transformer layers, and the embedding size is fixed at 512. We train COSMOS on A100 GPUs for 32 epochs, using a learning rate of $5\times 10^{-4}$ and a batch size of 1,024 for CC3M, or 4,096 for CC12M, YFCC15M and Merged-30M. Following DINO [7], we crop two global image views at scales between 0.4 and 1.0, resizing them to $224\times 224$ pixels, and six local image views at scales between 0.05 and 0.4, resized to $96\times 96$ pixels. Similarly, the two global text views consist of one to five sentences randomly sampled from the synthetic captions, while the six local text views consist of a single sentence sampled from the same caption set. The teacher momentum $\lambda$ is empirically set to 0.999 for CC3M and 0.99 for other datasets (more details in Sec. C.1).
@@ 4.1 Zero-Shot Image-Text Retrieval
As shown in Tab. 1, COSMOS achieves 53.1% and 40.1% R@1 score for image-to-text (I2T) and text-to-image (T2I) retrieval on MSCOCO [48]. These results are 12.3% and 10.3% higher than DreamLIP [91] (40.8% and 29.8%, respectively). When we switch to a larger pre-training set (CC12M), COSMOS further improves its performance, reaching 64.2% and 48.9% R@1 scores, outperforming CLIP [63] by 8.0% and 9.1%, respectively. When we employ YFCC15M [16] as the pre-training dataset, COSMOS attains a performance of 67.5% and 50.9% in terms of I2T and T2I R@1 scores. The highest performance of 68.0% and 52.5% on the MSCOCO [48] dataset is obtained by COSMOS, when Merged-30M is used for pre-training. COSMOS also achieves the best performance for image-text retrieval on the Flickr30k [84] dataset, achieving an R@1 score of 92.9% and 80.3% in terms of I2T and T2I retrieval when Merged-30M is used for pre-training, while SigLIP [88] attains only 89.9% and 73.3%, respectively. Overall, COSMOS outperforms all baselines on both datasets (Flickr30k [84] and MSCOCO [48]) in the image-text retrieval task.
@@
Table 3: Zero-shot semantic segmentation results in terms of mIoU (mean Intersection over Union) on the PASCAL VOC 2012 [22], PASCAL Context [55], Cityscapes [15], ADE20k [92], COCO-Stuff, and COCO-Object [5] datasets with and without background category. The vision encoder architecture is ViT-B/16. We also report the performance of our method enhanced with the SCLIP mechanism [77]. The best results are highlighted in bold. : Results copied from their work. : Results obtained using their official pre-trained weights.
@@ 4.2 Zero-Shot Image Classification
We observe in Tab. 2 that COSMOS outperforms the baselines CLIP [63], SigLIP [88], DreamLIP [91], MaskCLIP [21], and SF-CLIP [65] on most of the evaluation datasets with the same pre-training dataset. For instance, COSMOS pre-trained on CC3M reaches much higher accuracy of 37.1% on ImageNet compared to CLIP (23.9%) and SigLIP (25.9%). As expected, the classification task greatly benefits from the pre-training dataset size, i.e., COSMOS pre-trained on CC12M achieves 51.4% on ImageNet while on Merged-30M, it reaches 57.6%.
Additionally, COSMOS surpasses all baselines regardless of the pre-training set, in terms of average accuracy computed over the 11 datasets. For instance, when pre-trained on YFCC15M, COSMOS achieves an average accuracy of 50.0%, while SigLIP [88] and DreamLIP [91] reach 44.2% and 46.3%, respectively. It is worth noting that on two fine-grained datasets, i.e., Cars and SUN397, the accuracy improvements of COSMOS trained on Merged-30M is significant (67.8% vs 64.7% on SUN397, 31.2% vs 23.4% on Cars compared to DreamLIP).
Models trained on billions of samples, e.g., OpenCLIP [13], MetaCLIP [80], and Llip [41] achieve higher accuracy, implying that the larger the number of images the higher the accuracy for the classification task, since the model is required to learn the global information.